OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 3 7x - 5y = -11 Answer -3 -2 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system i got -2 ..

OpenStudy (anonymous):

3x - 6y = 3 7x - 5y = -11 this one we're going to need to transform one equation so when added one of the variable terms is eliminated. to do this i'm going to multiply 3x - 6y = 3 by -7: -7(3x - 6y) = -7(3) -21x + 42y = -21 and then i will multiply 7x - 5y = -11 by 3: 21x - 15y = -33 add them: 21x - 21x - 15y + 42y = -21 - 33 27y = -54 y = -2 B.

OpenStudy (anonymous):

b